use double angle formulas to find exact value:
SIN 660 degrees
COS 240 degrees

Answer by lwsshak3(11628)

use double angle formulas to find exact value:
SIN 660 degrees
COS 240 degrees
***
sin2x=2sinxcosx
sin660˚=sin(2*330)=2sin330cos330=2*-√3/2*1/2=-2√3/4=-√3/2
..
cos2x=cos^2(x)-sin^2(x)
cos240˚=cos^2(120)-sin^2(120)=1/4-3/4=-2/4=-1/2